larry is 8 years older than his sister. in 3 years, he will be twice as old as she is now. how old are they now?
Answer by mananth(16946)

Sister's age =x
Larry's age = x+8
After 3 years Larry's age = x+8+3=x+11
x+11=2x
2x-x=11
x=11, sister's age now
Larry = 11+8 =19 years